### v2

# Visualizing Quantum States: A Pilot Study on Problem Solving in Quantum Information Science Education

Abstract
In the rapidly evolving interdisciplinary field of quantum information science and technology, a big obstacle is the need to understand high-level mathematics to solve complex problems. Current findings in educational research suggest that incorporating visualizations in settings of problem solving can have beneficial effects on students’ performance and cognitive load compared to solely relying on symbolic problem solving content. Visualizations like the (dimensional) circle notation enable us to represent not only single-qubit but also complex multi-qubit states, entanglement, and quantum algorithms. In this pilot study, we aim to take a first step to identify in which contexts students benefit from the presentation of visualizations of single- and multi-qubit systems in addition to mathematical formalism. For this purpose, we propose test items and a complete methodology to assess students’ performance and cognitive load when solving problems. This is a pilot investigation with a large breadth of questions intended to generate hypotheses and guide larger-scale but more concrete studies in the future. Specifically, we compare two approaches: using the mathematical-symbolic Dirac Notation (DN) alone and using it accompanied by the (dimensional) circle notation. Analyzing time and performance, we find that most of the test items are appropriate for a heterogeneous target group as they can differentiate between the participants in regard to performance and time taken. In general, the A-B crossover structure of the study is suitable for investigating the benefits of visualization for problem solving with respect to length and feasibility. Future studies should, however, be more restricted in context because of the observable student and context dependence, with special interest lying in the further investigation of the Hadamard gate, the CNOT gate, and entanglement in multi-qubit systems.
I Introduction
The interdisciplinary nature of the field of Quantum Information Science and Technology ( Quantum Information Science & Technology (QIST)) necessitates endeavors to educate workforce with many different backgrounds [1]. This implies that the quantum education community needs to find ways to introduce the field of QIST to a broad audience. One way of facilitating such an introduction is the use of suitable representations, including visualization, to convey non-trivial quantum concepts.
Visualization is generally considered a useful tool in Science, Technology, Engineering & Mathematics (STEM) education [2, 3, 4, 5, 6, 7] and in QIST education in particular [8, 9, 10]. The field of QIST is based on the utilization of systems of multiple qubits, but the effectiveness of presenting visualizations of qubit states in problem solving, in terms of performance [11] and cognitive load [12], is currently unknown. Knowing under which conditions and in which contexts learners benefit from the depiction of visualizations, we would be able to offer QIST educators valuable advice for teaching by being able to recommend the most effective ways of utilizing visualizations. This is an important step towards solving the challenge of high heterogeneity of target groups in QIST education. It is unknown whether and under which conditions learning and problem solving are supported by providing visualization in addition to the mathematical symbolic representation (e.g., the DN) in QIST education. We identify three relevant questions in the context of visualizations in QIST education.
1.
In what contexts does incorporating visualizations support students’ performance in task solving on quantum operations and entanglement?
2.
How does the inclusion of visualizations affect cognitive load?
3.
How are performance and cognitive load affected by representational competence?
Currently, there are no validated multiple choice test instruments designed to answer these questions. In this study, we propose test instruments and a novel study design to answer these questions. We also present preliminary results, discussing the adequacy of the questions and conclude with important lessons for future studies. In this way, we take a first step in the investigation of whether visualizations provide a benefit to solve problems in QIST when provided in addition to the DN and under what circumstances such benefits occur. Further studies can build on test instruments, survey design, and preliminary results to answer the research questions above. In this way, the primary goal of this work is to generate hypotheses and identify promising avenues for further investigation, thereby laying the groundwork for more focused and robust studies on the effective integration of visualizations in quantum education.
The paper is structured as follows. First, we introduce the research background in quantum education and the visualizations used in II. We follow with a description of the study design in Section III and describe the creation of the test items in Section IV. The preliminary results are summarized, and the test instruments are discussed with respect to their suitability to the target group and indications of effects worth exploring further in Section V. We conclude and provide an outlook for future research in Section VI.
II Research background
The research background of this study includes the functions that multiple external representations (MERs) fulfill (Section II.1), an overview of educational research on representations in quantum education and visualizations of multi-qubit quantum states and entanglement, including the two representations investigated here, Circle Notation (CN) and Dimensional Circle Notation (DCN) (Section II.2), and the associated measures of interest, including cognitive load, multiple-choice test design and time taken (Section II.3).
II.1 Learning and Problem solving with visualizations in QIST
Based on educational research, it is well known that learning and problem solving in STEM can be supported by focusing not only on text-based and mathematical-symbolic representations (e.g., written text and formulas) but also on visualizations (e.g., pictures and diagrams) [13, 14]. From a theoretical perspective, the benefit of learning with text accompanied by visualization can be explained by efficient information processing and the dual-channel structure of the working memory. When learning with Dirac Notation alone, the processing load of information is solely on the verbal channel of working memory. However, when learning with text and visualization, both the verbal and visual channels of working memory are activated and integrated in the construction of coherent mental schemata [15, 16]. This distributed information processing across two channels reduces the load on each single channel and, thus, the risk of cognitive overload.
Within her Design, Functions, Tasks (DeFT) framework, Ainsworth [17] has formulated three main functions of MERs to support learning. First, MERs can complement each other by containing different information or supporting different processes. Second, they can constrain each other, e.g., by familiarity of learners with a subset of the MERs or inherent properties. Third, incorporating MERs can construct deeper understanding as learners are confronted with the abstraction of the underlying knowledge structures, the extension of knowledge to an unknown representation or improving the understanding of the relations between different representations.
However, incorporating additional representations can also impose additional cognitive load on students, as they must understand not only how each representation conveys information but also how to translate between them. Therefore, the learning effectiveness of MERs does not only rely on the learning material but also on the characteristics of the learners [17, 18]. To learn and solve problems with MERs efficiently, students must possess representational competence [19]. Rau divides representational competence with multiple visual representations into conceptual competencies (visual understanding, the ability of connecting the representation to concepts and understanding of the main features, and connectional understanding, the ability to connect and compare the representation to other representations), perceptual competencies (visual fluency, i.e., efficiency of connecting the representation to concepts and the ability to work effectively in the representation, and connectional fluency, i.e., efficiency and flexibility of connecting multiple representations to each other and switching between them), and lastly meta-representational skills (i.e., the ability to choose suitable representations) [5].
II.2 The Multiple Representations in QIST education
Research was carried out on learning quantum physics with representations such as the Dirac and vector/matrices notation of single-qubit states and operations, considering meta-representational competencies like assessing the suitability of the two representations [20]. The authors found that alternatively to investigating the learning difficulties of students, one can find out how representations are used to solve tasks in order to find more productive teaching approaches. Extending this work to the utilization of visualization techniques would be a natural way of moving forward.
Visualizations are widely used in teaching and learning quantum physics, and some of them have been investigated regarding their efficacy for learning and corresponding learning difficulties. Some basic visualizations of qubit states, such as the unit circle, the Bloch sphere, coin analogies, and quantum circuits, are summarized in [21]. Different visualizations of a beam splitter were tested [22] which later led to the creation of an interactive visualization tool of the Mach-Zehnder interferometer [23], which can be see as an optical implementation of a single-qubit system. Student difficulties with the Mach-Zehnder interferometer when learning with an interactive visualization were also analyzed in [24]. The Bloch sphere is a popular representation for single-qubit systems that can help make respective quantum operations more intuitive by relating them to rotations and projections in three-dimensional space. Difficulties with the Bloch sphere, the concept of a basis and the measurement process were analyzed in [25]. They found that their tutorial with the Bloch sphere increased performance in regard to measurement probabilities and outcomes, and that the Bloch sphere helps identifying states that are equivalent, up to a global phase. In [26], a new learning sequence with the Mach-Zehnder interferometer has been proposed, incorporating the unit sphere to visualize polarization, an interactive GeoGebra animation, a graph experimental setup, and more abstract illustrations and mathematical formulas. Ideas such as these highlight the diversity of representations even in such a restricted context as the Mach-Zehnder interferometer. In the following, we briefly describe existing visualization methods of multi-qubit systems and the motivation to study their impact on task solving ability.
Representations of multi-qubit system
QIST is based on mathematics, in particular, linear algebra in complex Hilbert spaces [27, 28]. This is because quantum states can be represented as vectors in these Hilbert spaces, and quantum operations correspond to unitary operations or projections. When moving to multi-qubit systems, understanding this mathematical picture becomes much more challenging. To combat the high level of abstraction in this mathematical field and gain some intuition, a multitude of visualizations are available. More advanced visualizations include broad geometric depictions of quantum states and entanglement as shown in [29, 30, 27, 31]. It can also be used to show multi-qubit states [32, 33, 34, 35]. The diagrammatic ZX-calculus [36, 37, 38, 39] can be used to describe specific quantum operations and whole algorithms in a “LEGO-like” manner [40], simplifying the standard quantum circuit notation. With so-called color code, quantum error correction algorithms can be visually depicted as acting on hypercubes or hypercube-like lattices [41, 42, 43, 44]. The DROPS representation utilizes generalized Wigner functions to visualize quantum states, entanglement, and quantum operations in an operator-basis [45, 46, 47].
All of these visualizations use other bases than the computational (0,1)-basis. However, in this work, we focus on the computational basis as it is often used for introductory purposes [48]. A dynamic visualization of quantum states and circuits was created in [49], based on the diagrams of state visualization [50], resulting in a quantum puzzle game [51]. The Circle Notation is a visual representation of the complex amplitudes of states in the computational basis using circles with gauges [52]. While in CN, amplitudes are aligned in a row as you would when writing states in the Dirac Notation, qubits can also be assigned axes in space [53, 54]. This dimensional approach has the possible benefit of making quantum operations more intuitive and can be combined with CN (then called DCN). It can be used to visualize the entanglement properties of qubit systems [55].
Currently, we do not know of results showcasing the efficacy of different visualizations of multi-qubit systems regarding learning outcomes and task solving ability. As similar processes are underlying these two [11] and in a similar line of thought to [20], we believe that gaining insight into task-solving abilities and processes with different visualizations can help improve teaching and learning by answering when and how best to use which visualization techniques. An increase in task solving ability connected to a better understanding of the processes could also lead to a better understanding of the underlying concepts of superposition and entanglement, which have been shown to have a high intrinsic susceptibility to misunderstandings [24, 56]. In the following, we further introduce and discuss the visualizations analyzed here, the CN and DCN.
Circle notation
The CN uses filled circles with pointers to visualize the amplitudes of a quantum state . We refer to qubit states using the computational basis, which means that is written in binary form, with the rightmost digit being the least significant and also the number corresponding to the first qubit. The magnitude is represented by the radius of the filling circle and the gauge at an angle , measured counterclockwise from the vertical axis, as shown in Figure 1. Because a quantum state is normalized, that is, , the sum of the areas of the inner circles is . Each area can be interpreted as the Measurement Prob. of the corresponding basis state.
As shown in [52], CN can be used to visualize quantum algorithms ranging from single-qubit operations over three-qubit quantum teleportation and swap tests, to four qubit amplitude amplification used in Grover’s algorithm and even eight-qubit quantum phase estimation used in Shor’s algorithm involving 256 circles. Beyond these numbers of qubits, the visualization gets unfeasible as the number of basis states is doubled with each additional qubit. This is a major limitation of the visualization. However, most of the known quantum algorithms can be implemented on a basic level using eight qubits, and therefore, CN provides a useful visualization of quantum algorithms in general.
Dimensional circle notation
In dimensional notations, every qubit is assigned an axis in space [53, 55]. DCN is such an extension of CN. For comparison, Figure 2 shows the same state as Figure 1 in DCN. For further clarification, we also compare the use of a Hadamard gate that creates or destroys superposition for CN and DCN in Figure 3. In principle, DCN can be used to visualize the same problems as CN. When visualizing systems of more than three qubits, multiple cubes or hypercubes can be arranged in lattices. Also, the visualization can be modularized as a mixture of the one-dimensional CN and the dimensional approach [55]. It becomes more difficult to implement in practice when trying to visualize systems of more than six qubits, because the visualization is less compact than the CN. It also leads to the necessity of page scrolling if five three-qubit systems are visualized in a given question, using standard screen sizes.
Entanglement in the computational basis
In the computational basis, separability of pure states can be seen as symmetry apart from a complex factor with respect to every qubit of the system [55]. This is described by the following theorem [57, 55]:
Theorem 1
Let . An -qubit state is 2-2n-1 separable into if and only if for all either or there exists a ratio such that .
This ratio is the complex factor that can be seen in separable states in dimensional notations. To see this factor in two-qubit systems, one needs to observe whether there is symmetry along the axis of one qubit, apart from a complex factor, the ratio . In the absence of symmetry, the system is entangled. Similarly, in three-qubit systems, one can spot whether a qubit is separable from the rest of the system, i.e., not entangled with it, if there is symmetry in regard to the plane perpendicular to that qubit’s axis. A system is called partially separable if only a subset of the qubits of that system are separable from that system. This is shown in Figure 4. In systems of more qubits, this theorem can be generalized to --separability and the separability from a two-qubit system to the rest of of the system can be spotted, albeit it is much more complicated to do so as it involves the comparison of ratios along the three axes of a two-qubit state, including the diagonal [55].
Referring to the DeFT framework, utilizing visualizations such as CN and DCN to support problem solving makes use of the representations complementary functions, by visualizing qubit characteristics and gate operations in addition to the common mathematical-symbolic description. As students are provided with both representations, they have further access to QIST content knowledge as compared to the mathematical description alone. In addition to this, their use facilitates the understanding of the corresponding mathematical descriptions. When using both visual and mathematical representations, students can be supported in extending existing knowledge structures to mathematical descriptions. For example, in the context of entanglement, dimensional visualizations like DCN provide new representations of the separability of qubit systems or even subsystems [55].
Like any representation, there are shortcomings that are not easily alleviated. Firstly, in one-qubit systems, the Bloch sphere and the generalization to multi-qubit systems in form of the QuBeads representation [58] might make operations and entanglement more intuitive by relating them to rotations in space, and could therefore be used to complement mathematical notation more effectively. Secondly, these representations take up a lot of space and can display information only approximately, making mathematical symbolism and the circuit notation that can give an overview of a complete quantum algorithm the more concise choice to constrain interpretation. Lastly, only numbers are visualized, while algorithms should work on arbitrary quantum states and, therefore, their function can only be proven using variables. The additional visualization can only add an intuitive understanding of the underlying workings of quantum algorithms and perhaps deepen it. Due to these shortcomings, we recommend usage in teaching and learning in addition to mathematical symbolism and possibly other representations, instead of using solely visualization, in order to complement, constrain interpretation, and construct deeper understanding, according to the functions of the DeFT-framework.
II.3 Relevant outcome measures and test construction
Cognitive Load
It is essential to consider the cognitive load imposed on students when designing tasks and learning materials, as working memory capacity is a major limiting factor in learning and problem solving [12]. In the original version of the Cognitive Load Theory (CLT) Sweller divided cognitive load into three categories: extraneous cognitive load ( Extraneous Cognitive Load (ECL)), intrinsic cognitive load ( Intrinsic Cognitive Load (ICL)) and germane cognitive load ( Germane Cognitive Load (GCL)) [59]. ECL is cognitive load imposed on the learner by design, i.e., how the material is presented. By reducing ECL there is more cognitive capacity for goal-relevant cognitive processes. ICL, on the other hand, is cognitive load inherent to the task or the learning content and can only be reduced if the task is made simpler. This could be something that MERs can accomplish by providing access to alternative strategies. GCL is the effort to construct cognitive schemata, that is, mental patterns and categorizations that allow students to infer meaning from the learning material. To support students in the construction of cognitive schemata, materials should be designed such that ECL is low, but GCL is high.
When incorporating a visualization in addition to the DN, learner characteristics can determine the imposed cognitive load and whether the visualization helps or not. For example, a novice might, in some cases, benefit from using the visual representation of a complex number in CN, while an expert might never need to use a second representation, and the additional representation just adds to ECL. This effect is called the expertise reversal effect [60]. On the other hand, someone who understands the action of gates and the definition of separability in dimensional notation, i.e., has representational understanding, could benefit from DCN in these contexts. More generally, these two effects are forms of aptitude-treatment interactions [61].
Multiple-choice test item design
The test items for this pilot study were chosen as multiple choice because of the several advantages they have over open questions [62]. Multiple choice items are much more economical, allowing for faster score evaluation. If validated in detail, the scoring becomes accurate, as open-ended questions would leave room for interpretation. In addition, multiple choice items allow for assessment of time taken without having to control for abilities of students like how fast they can write / type that are probably not relevant to the research questions. Multiple choice tests also take advantage of the testing effect, enhancing the recall ability of students [63]. Using eye tracking in future studies, thought processes during task solving can be analyzed, further mitigating the advantage of open questions in providing insight into thought processes.
When designing multiple choice items, it is important to also consider distractors as part of the item [64]. Their design should be competitive with the correct answer, that is, they should be plausible alternatives [63]. It should be considered how participants might think when the item is presented to them and common errors. For example, common errors or misconceptions in the context of this study can be confusion of the order of qubits, measurement probability as just magnitude and not magnitude squared [65], confusion of basis states with qubits [66], or confusing separability with entanglement [67]. If distractors do not follow common misconceptions, they should share characteristics with the correct answer or the other distractors in order to be competitive [68]. Distractors should be designed in a way that leaves no hint towards the correct answer without reading the question.
Time taken
The time it takes a student to answer a question correctly depends on different factors. The distance-difficulty hypothesis proposed in [69], and revised in [70], predicts that the logarithm of the time taken to answer the questions is proportional to the distance of the students’ ability to the difficulty of the question. In [71], it was confirmed that in complex tasks, students who took longer also performed better. In [72], it was found that children with higher levels of ability take longer to incorrectly solve the questions, so the distance-difficulty hypothesis holds in this context. Considering the impact of providing a visualization on the difficulty of the question, if one finds that participants take longer times on questions with a visualization, it could mean that the visualization provided complementary strategies to solving the question that brought the difficulty of the question closer to the level of ability of the student.
When only considering time taken for correctly solved questions, most of the results where the question was above the skill level of the participant are eliminated (except for those questions that were correctly solved by random chance). Therefore, when considering time taken only for correctly solved questions, one considers more of the students that are above or at matching ability level to the question’s difficulty. If a question becomes trivial for a student when presented with a visualization, they will take less time for correctly solved questions. Therefore, when requiring that both the question without visualization and with visualization be solved correctly, on average the time taken could be lower if the question did, in fact, get easier by providing a visualization.
The time taken for task solving is a multidimensional construct. Clear results are expected to be very difficult to obtain, as there are many factors that play a role. However, according to the distance-difficulty hypothesis [69, 70], the time taken could be used as a metric (albeit weak) to determine how suitable the given tasks were to the level of ability of the students. Here, all answers, correct or incorrect, should be considered. In addition, a close analysis of the dependence of the difficulty of the question on the time taken for correct answers could provide further details on the benefits of providing a visualization.
III Research Questions
Based on current educational research, we expect that the use of visualizations such as CN and DCN can utilize the known advantages of learning with MERs in QIST by constructing deeper understanding of entanglement and complementing the mathematical symbolic DN by offering new strategies to predict the outcome of quantum gate operations. In this study, our objective is to take the first step of the design and validation of test items that can be used to verify this expectation by comparing the task solving ability of participants in DN with those where DN is accompanied by CN or DCN. Such tests should take into account the representational competence of the participants. This competence is divided into conceptual competencies related to visual and connectional understanding and the perceived cognitive load during tasks without and with visualization [5]. Here, we refer to representational understanding instead of visual understanding, including the competency with complex numbers in mathematical symbolic notation. It is not clear whether this is an adequate generalization, that is, whether Rau’s construct of representational competence adapted to incorporating representations beyond those that are purely graphical still predicts performance increases. It is also unclear whether this still holds in task solving rather than learning, as the two differ in the decision-making phase [73]. Because the supposed benefits, in addition to being likely learner-dependent, are also likely to be context-dependent, we explore the effects of presenting a visualization in different contexts, that is, measurements, various unitary operations, and entanglement & separability.
The aim of this pilot study is to create test items and a suitable study design to answer the following overarching questions, and taking a first step of examining the validity of these items.
1.
In what contexts does incorporating visualizations support students’ task solving ability on quantum operations and entanglement?
2.
How does the inclusion of visualizations affect cognitive load?
3.
How are performance and cognitive load affected by representational competence?
Therefore, we formulate the following research questions.
- RQ1
Are the proposed test items suitable in terms of difficulty for a heterogeneous target group?
- RQ2
Which test items should be omitted or redesigned?
- RQ3
Is the study design suitable for answering the overarching questions?
In addition, we intend to find indications of effects in regard to the overarching questions that are worth exploring further in future studies.
IV Methods
In order to gain insights into the effects and relevant conditions of including visualizations like CN and DCN in students’ problem solving, we decided to compare the performance and cognitive load of participants with and without visualization intrapersonally and relate them to representational competence. As is the nature of a pilot study, our aim is to learn about the suitability of the test items that are presented to the participants and the study design. All test items were reviewed by experts using the design principles described in Section II.3. They are provided in the supplementary material, the correct answer being highlighted. The order of the answers was randomly assigned to each participant. In Section IV.1, we define the overall structure of the pilot study. We then describe the group of participants in IV.2. In Section IV.3, we describe the introductory slides used. In Sections IV.4 and IV.5, we show the design of the questions on representational and connectional understanding. In Section IV.6 we describe the test items of the main parts of the survey. We finish with the cognitive load test items in Section IV.7.
IV.1 Structure of the study
The study consisted of three online multiple choice tests about one-, two-, and three-qubit systems, respectively. Each of the tests took about 45 minutes. The participants received 10€ compensation per test.
In each survey except for the first, the participants were randomly assigned to two groups, where one group was presented with CN and the other group was presented with DCN. Each survey was separated into two parts, A and B, where questions using the same contexts (gate operations, measurements, entanglement, and separability) were asked in part A and B. One group was presented with visualization in part A and without visualization in part B (group vis-math), while the conditions were reversed for the other group (group math-vis). After each part, we measured the perceived cognitive load. Before the surveys, we showed introductory slides to the respective visualization, operations and entanglement, and asked questions measuring representational competence. In the end, we asked for demographic information (age, level of educational achievement and field of study). The whole structure of the surveys is shown in Figure 5.
IV.2 Participants
In the survey of the one-, two- and three-qubit systems, there were 17, 21 and 15 participants. At the end of the first survey we asked for demographic information.
Of the 17 participants in the first survey, eight stated that they had not been introduced to the DN before and nine participants stated that they were introduced to CN before the survey. From the participant group for the study on one-(two-)(three-)qubit systems, 1(1)(1) was in the age range 18-21, 12(11)(9) were in the age range 22-25 and 3(2)(3) were in the age range 26-30. Furthermore, 16(14)(13) were male participants and 1(1)(1) female. The level of educational achievement was high school graduation for 4(3)(4) participants, completed Bachelor’s degree for 9(7)(5) participants, completed master’s degree for 3(3)(3) participants, and completed Ph.D. for 1(1)(1) participant. There was 1(1)(1) participant in the field of Economics, 8(8)(7) participants in Information/Computer Science, 3(3)(2) participants in Computer Science and Engineering, 4(4)(4) participants in Mathematics, and 1(1)(1) participant in Physics.
The small sample size of this pilot study does not allow a reliable quantitative analysis. Therefore, we perform a qualitative analysis on the indications given by the data analysis and present the detailed quantitative results in Appendix A. In the main body of this work, we describe the suitability of the questions for the participants, where it can be inferred from the answer correctness, time taken, and cognitive load. Depending on whether the average correctness was above 84%, between 51% and %84, and below %51, we label each question type of the main parts of the study easy, moderate, or hard and with low variance or high variance, depending on how large the differences between groups, with or without visualization were. Larger differences would mean either higher dependency on learner prerequisites or the presentation of visualization that are of value to explore in future studies.
IV.3 Introductory slides
At the beginning of each survey, participants were given the following introductory slides (in the corresponding visualization, CN or DCN, that they were assigned):
Visualization of qubit states
Action of single-qubit gates in single-qubit systems
Action of multi-qubit gates in multi-qubit systems (only shown in the second and third survey)
Measurement probabilities and the resulting state after measurement
Separability and Entanglement (only shown in the second and third survey)
An example slide is shown in Figure 6. The introductory slides of the one-, two-, and three-qubit survey are shown in the supplementary material. They were designed in a way to introduce the visualization complementary to the DN, and to show the strategies unique to each representation. The introductory slides are supposed to reduce the influence of previous knowledge on the test results. This was done because the main purpose of the questions is to find out whether students benefit from the visualization, previous knowledge being unaccounted for, due to a lack of test instruments for this construct.
IV.4 Representational understanding
In the one-qubit survey, participants were asked one question on the multiplication of complex numbers and one on the Euler formula (for example translating between and ), and similar questions using only CN. For this, four different questions were designed, two of which each participant received in CN and two of which each participant received in mathematical notation, randomized. An example is shown in Figure 7. These questions are limited in their applicability to basic arithmetics in DN and CN and do not capture spatial reasoning ability that is supposedly necessary for a benefit of DCN in multi-qubit systems. These questions were therefore only used in the one-qubit survey. All representational understanding questions are shown in the supplementary material. The questions were designed to include possible misconceptions about CN: students might think that the gauge pointing to the right indicates a phase of 0, and that adding a positive phase means that the gauge is turned counterclockwise.
Participants were asked their confidence on each of these questions on a Likert scale from 0 (random guess) to 4 (very sure). The responses were evaluated giving participants a weighted score as proposed in, e.g., [74] as having the benefit of reducing statistical errors, but also discussed as possibly influenced by characteristics of the participants such as self-esteem. The usage of the visualization is a deciding factor for a benefit of the visualization, if there is one. We therefore expect the weighted measure to act as a reliable moderator. We did not repeat the representational understanding test in the other surveys due to time restrictions.
IV.5 Connectional understanding
To measure connectional understanding, we measure the translation competence of participants between DN and visualization. After the introductory slides, the participants were asked four questions (six questions in the two-qubit survey) to translate between their assigned visualization and the DN. Two (three in the two-qubit survey) of these questions were about translating from DN to visualization, and two (three) were about translating back. One question was about translating phase, one was about translating magnitude, and in the two-qubit survey, we added one question that asked about both. An example task in the three-qubit survey is shown in Figure 8. All translation competence questions of the one-, two-, and three-qubit survey are shown in the supplementary material.
The questions measuring connectional understanding were designed in a way to check either the participants’ ability to translate phases or to check the participants’ ability to translate magnitudes. The distractors were chosen such that no answer stands out regarding a specific aspect, or all stand out, i.e., so that participants can not infer hints towards the correct answer without looking at the question. In addition, the questions are rather diverse, showing different types of combinations of amplitudes.
IV.6 Quantum Operations & entanglement
We tested the visualizations in one-, two-, and three-qubit systems. The X and Hadamard gates were tested for all sizes of qubit systems and the Z gate in one- and three-qubit systems. In addition, the measurement probabilities and the resulting state after measurements were asked in all surveys. The phase gate was only tested in the one-qubit survey, while in the two- and three-qubit surveys the CNOT gate, (partial) separability, and (partial) entanglement were also tested. Figure 9 shows an example task of the effect of a Hadamard gate in a two-qubit system. A short explanation of how a given operation acts / how entanglement or separability is defined was given to reduce the relevance of previous knowledge with the same reasoning as showing introductory slides. In addition, doing so reduces the possibility of a false interpretation of what is asked, according to the recommendation in [68].
Questions of the one-qubit survey
In the one-qubit survey, we first asked about the equivalency of states that only differ by a global phase to see whether it would be easier to see these phase differences in the CN representation. The distractors were chosen to include confusion with shifting only one of the two phases and swapping only the magnitudes. The -gate questions in one-qubit systems were designed to include the confusion about which part of the amplitude to swap, i.e., one distractor was chosen such that only the magnitude is swapped and one such that only the phase is swapped. The last distractor was chosen as equalizing the magnitudes. For this question, participants needed to refer to the equivalency of two states that differ only by a global phase (). The -gate question in one-qubit systems was designed to be more challenging, as a mixture of destruction and creation of a superposition needed to be applied. It was meant to involve the complex number distributive multiplication competence, but we found out in discussions with the participants that this particular question was too complicated, and we diverted to only asking about situations where amplitudes are either created or destroyed, but not both at the same time, in the two- and three-qubit surveys. Part of the complication was that, similarly to the -gate question, the correct answer is shifted by a global phase (). These global phase differences might be an important disadvantage of CN compared to the Bloch sphere, where the global phase information is left out by construction. The -gate question was chosen to contain confusion of a phase shift with a phase shift, the misconception that -gates might also swap magnitudes and a combination of both. In the phase-gate questions, we added distractors of having the phase shift in the wrong direction, or doubling the phase shift in either direction.
With the measurement questions, we intended to find out whether the depiction of measurement probabilities by the area of the circle helped to make statements about these probabilities. We included the following distractors: only taking the real value without squaring it, not being able to make any statement about the measurement, and the measurement result always being 1. The other measurement question was about the resulting state after measurement and included states that were not normalized and states where nothing has changed.
Quantum Operations in two- and three-qubit systems
The questions of the two- and three-qubit surveys were designed similar to the one-qubit survey, although we omitted the phase gate and Global Phase Equiv. questions in favor of entanglement and separability. When designing these questions, distractors were included in which the gate was applied to a wrong qubit or to both qubits (in the case of the CNOT question, we decided on a SWAP gate as a distractor). As a last distractor, a state with similar features to the others was chosen such that no state stands out regarding a specific feature. In some cases where two states would be the same using these approaches, we added two such states as distractors.
The Guess-the-Gates (GtG) questions of the two-qubit survey, where it was asked which gate combination was applied, were designed to include all combinations of or gates, where . The GtG questions of the three-qubit survey were designed similarly, but including the or operations in part A and measurement operations in part B.
The Measurement Prob. questions were designed similarly to those of the one-qubit survey, except that we referred to percentage differences in measurement probabilities in the two-qubit survey and reverted back to multiples in the three-qubit survey. We expected that the visualization would not provide much aid here, as it could be difficult for students to add all the areas together. However, it might provide hints towards the correct answer for students that are not as comfortable with the fractions in mathematical symbolism. The distractors show normalized states in which the measured qubit is always in the measured state, 0 or 1. In this way, the Measurement State questions were designed to include fewer misconceptions. In one Measurement State question, we asked participants to identify the correct phases after measurement. Here, a global phase was left constant and four answer possibilities were included, the correct being (, ), and the incorrect (, ), (, ), and (, ). In the amplitude Measurement State questions, the amplitudes were swapped around.
Entanglement and Separability
For each of the two- and three qubit survey, we designed four questions on the topic of entanglement. In the two-qubit survey, we asked two questions about finding a state that is separable and two questions on finding a state that is entangled. One of these questions asked about magnitude entanglement with all-equal phases, the other about phase entanglement with all-equal amplitudes. For the magnitude Entanglement questions, amplitudes were switched around such that only one state is entangled / separable. In the phase separability questions, cases where the ratios (see Section II.2) where and , and , or and were added as distractors. Note that there exist many more entangled states than separable states [75]. For this reason, the question find the separable state has many more choices regarding the distractors than the question find the entangled state. In the latter case, the choices were restricted to equal ratios. We chose a state with and as the entangled state, and varied the direction and value of the phase differences, as well as the phase of the state to design the other distractors.
In the three-qubit survey, we asked about fully entangled or fully separable states and showed only partially entangled states as distractors. To design the question to identify entangled states, we started with one fully phase or amplitude entangled state, and constructed similar states that were symmetric in terms of one qubit, making that qubit separable from the system. To design questions asking to find fully separable states, we started with a fully symmetric state as the correct answer and applied CNOTxy and CNOTxz gates for two , as well as a mixture of the answer possibilities, to arrive at the distractors. The amplitude questions only had four amplitudes that were non-zero, in the phase questions all amplitudes were non-zero. All magnitudes were chosen to be equal in all Separability and Entanglement questions to reduce complexity.
Similarly to the Measurement phase and amplitude questions, the number of questions in order to include all combinations of entanglement-phase, entanglement-amplitude, separability-phase, and separability-amplitude, with and without visualization, should be doubled. Again, this would only be feasible in a more focused study.
IV.7 Cognitive load
The perceived cognitive load was measured after each part of the survey, A and B. The following Cognitive Load test items were chosen, as suggested in [76]:
- ICL question 1
For the tasks of survey part A/B, many things needed to be kept in mind simultaneously.
- ICL question 2
The tasks of survey part A/B were very complex.
- GCL question 1
For the tasks of survey part A/B, I had to highly engage myself.
- GCL question 2
For the tasks of survey part A/B, I had to think intensively what things meant.
- ECL question 1
During the tasks of survey part A/B, it was exhausting to find the important information.
- ECL question 2
The design of the tasks of survey part A/B was very inconvenient for learning.
- ECL question 3
During the tasks of survey part A/B, it was difficult to recognize and link the crucial information.
Possible answers were on a Likert scale from 0 (completely wrong) to 6 (absolutely right).
The naive questionnaire of [76] was chosen to measure GCL, because we concluded that the GCL test items of the first iteration fit our use case of problem solving tasks better. Due to the limited effects of the naive GCL items in [76], the GCL measurement should be critically viewed. However, the items used ask about the perceived mental effort of the participants (the perceived level of cognitive engagement and the intensity of thinking), which is supposedly linked to GCL [77].
V Results & Discussion
In the following, we refer to group math-vis as the group that was assigned questions without visualization in survey part A and questions with visualization in part B. Vice versa, the group vis-math was assigned questions with visualization in survey part A and questions without visualization in part B.
Due to the low number of participants, statistical analysis is not feasible, and we refrain to descriptive statements. The quantitative data is shown in the Appendix A.
We discuss the results obtained regarding answer correctness and time taken for correct answers for the one-qubit (Section V.1), two-qubit (Section V.2.1 and V.2.2) and three-qubit (Section V.3.1 and V.3.2) surveys with the intention of learning about the suitability of the questions and indications of effects to explore in future studies. As described in [78], the difficulty of an ideal item with three distractors is 0.74. We deem an item difficult at or below an average correctness of 0.5, moderate between 0.51 and 0.84, and easy at or above 0.85 [79]. We also examine the variance of the results for each type of question with and without visualization and the differences between two groups to separate the questions into high or low variance. We consider questions with moderate difficulty and high variance to be of the highest value for further exploration. We provide such an overview of the discussion in Section V.4. A more detailed analysis can be found in Appendix A.
V.1 One-qubit systems
In the first survey, participants were asked to find
a state that is equal to another state apart from a global phase
the resulting state after application of an X gate (only in part A) or a Z gate (only in part B), phase gate, and a Hadamard (H) gate
the probability of finding a particular result upon measurement of a given state, and
the resulting state after measurement.
The average scores were without visualization and with visualization, at average times of and seconds, respectively. In general, the questions seemed suitable for the participants, as the mean scores were mostly above 50%, except for the Measurement Probability question. The average scores for single-qubit questions are shown in Table 4 and the average times for correctly solved questions are shown in Table 5. Figure 20 shows the combined average score per question and Figure 21 shows the combined average time taken (for all answers, not only the correct ones).
Representational understanding
Participants were asked to calculate Euler formula identities and similar questions to translate between CN and a complex number. In addition, they were asked to multiply the numbers written in DN and CN as shown in Figure 7. Then they were asked to rate their confidence in the question from 0 (random guess), 1 (very unsure) to 4 (very sure). The correct answers were weighted (multiplied by , where is the confidence). Detailed results are shown in Section A.1.1.
None of the participants scored higher with CN than with DN (weighted or unweighted), suggesting that the representational understanding of most participants of the visualization was lower than that of DN. It is possible that presenting the visualization to students with high representational understanding with CN but low representational understanding with the DN has greater benefits.
All in all, we see that possible benefits of visualization are learner- and context-dependent and, if it is, probably not strongly dependent on the presence of the visualization, that is, we expect small effects if any. It is worth investigating further under which circumstances benefits occur. In one-qubit systems, the CN is merely a visual representation of complex numbers, providing an aid, in theory, complementary to the DN. In two- and three-qubit systems, DCN is an extension of CN that could provide additional benefits due to the introduction of dimensionality.
Translation competence and cognitive load
On average translation competence, 12 students achieved perfect scores and 5 students achieved a score of 0.5. The students who did not score perfectly also scored lower in the rest of the survey (0.50.21 to 0.750.21 without visualization and 0.630.19 to 0.690.25 with visualization) and took a smaller amount of time on the correctly solved questions (8274 to 11471 seconds without visualization and 8657 to 8955 seconds with visualization) than the students who solved the translation questions perfectly. Average ECL (2.881.67 with visualization and 2.761.54 without) was lower than the average ICL (3.750.97 with visualization and 3.641.04 without) and GCL (3.941.26 with visualization and 4.131.41 without). In the difference of perceived ECL of survey parts supported by visualization minus parts not supported by visualization, the numbers fluctuate much more (as can be seen in the high variance, depicted in Table 2). This high variance could indicate a large dependency of ECL reduction due to visualization on learner characteristics.
V.2 Two-qubit systems
In both parts of the second survey, participants were split into two groups, one presented with CN and one presented with DCN, and were asked to find
the resulting state after application of an X gate and a Hadamard (H) gate,
the probability of finding a particular result upon measurement of a given state,
the resulting state after application of a CNOT gate,
a separable state and an entangled state (between non-separable, or non-entangled states, respectively), and
a combination of two gates that were applied to transform one given state into another given state (“Guess-the-Gates” (Guess-the-Gates) questions).
The average scores for two-qubit questions are shown in table 6 and the average times for correctly solved questions are shown in table 7. In the latter, the combined score of part A and part B show only cases where participants managed to solve both questions, i.e. both questions on the same concept with and without visualization, correctly. Figure 20 shows the combined average score per question and Figure 21 shows the combined average time taken (for all answers, not only the correct ones).
V.2.1 Two-qubit systems with Circle Notation
In the two-qubit survey with CN, the average scores were without visualization and with visualization, at average times of and seconds, respectively. In both survey parts, we see group math-vis scoring higher or the same as group vis-math on all questions except the Separability and Entanglement questions, where group vis-math scored higher. This could be explained with learner characteristics, such as the ability to identify entanglement by visual means or using mathematical strategies. The vis-math group scored higher on these questions when presented with visualization. The H gate question shows a large difference in score between the two groups, where group math-vis scored 0.8 without visualization and group vis-math only 0.2. This could hint towards the Hadamard gate question performance being especially prone to learner prerequisites.
When presented with the visualization, participants took longer when solving questions correctly, at seconds seconds. This could be the result of alternative strategies to solve the question that the visualization provides, which bring the question’s difficulty closer to the level of the participant’s ability, according to the distance-difficulty hypothesis [69, 70].
Translation competence and cognitive load
Similarly to the one-qubit survey, participants scored in translating between CN and DN. Again, we see that participants who scored perfectly when solving these questions also scored higher in the rest of the survey ( without and with visualization) than participants without a perfect score ( without visualization and with visualization).
The group of participants with perfect translation scores also took more “reasonable” amounts of time to correctly solve questions, that is, the variance in time taken was lower (1.670.65 seconds without visualization and 2.030.67 with visualization) than in the group without a perfect translation competence score (1.130.79 seconds without visualization and 1.781.52 with visualization). Presumably, students that don’t know the answer and choose to take a guess will take a short amount of time, while students that try hard to find an answer in tasks where this would not be necessary for students with high previous knowledge, will take a long time. Therefore, a higher variance in the time taken for task solving could be expected for lower-performing students [69, 70].
The results are also similar to the results of the one-qubit survey with respect to perceived cognitive load. Average ICL ( without and with visualization) and GCL ( without and with visualization) are higher than average ECL ( without and with visualization). It seems that on average, the questions succeed in offering a high degree of intrinsic difficulty, as indicated by the comparatively high ICL.
V.2.2 Two-qubit systems with Dimensional Circle Notation
In the two-qubit survey with DCN, the average scores were without visualization and with visualization, at average times of and seconds, respectively. In all question types, one group performed better than the other group in both part A and part B (regardless of whether visualization is presented). This consistency can possibly be explained by learner prerequisites predicting the performance in these questions better than whether visualization is presented or not. In addition, this could be a manifestation of the context dependency of student competence, i.e., a student being competent with one quantum process does not necessarily mean that the student is also competent with other processes. The differences between the scores of the groups is particularly apparent in the H gate, entanglement, and CNOT gate questions in favor of group vis-math (0.5 to 0.17, 0.7 to 0.42, and 0.9 to 0.69, respectively) and the Guess-the-Gates and separability questions in favor of group math-vis (0.42 to 0.1 and 0.67 to 0.5, respectively).
The H gate and Guess-the-Gates questions seemed quite difficult, as the average scores in these questions were below 0.4. Feedback from some students after the test suggested that the GtG questions were particularly challenging and exhausting.
Considering the amount of time spent for correctly solved questions, we see outliers of more than 6 minutes in the group vis-math on the Z gate A, H gate A, and H gate questions of part B. Participants took, on average, longer when presented with visualization than when they were not. This is largely due to these outliers, as, when removing the outliers, the difference is only marginal (1 minute 58 seconds without visualization to 2 minutes and 5 seconds with visualization).
Translation competence and cognitive load
In the two-qubit survey, translation competence was lower in the DCN group at than in the CN group at . In addition, no one in the DCN group correctly answered all the translation questions. The average scores in this survey were also lower than in all other surveys, at without visualization and with visualization.
For participants of the DCN group, ICL was slightly higher than for participants of the CN group when solving math questions at , as compared to , and we see indications of reductions in ICL and ECL when participants are presented with DCN of and , respectively, with a comparatively large fluctuation of this difference. We do not see a similar reduction in GCL as we do in the CN group. The cognitive load results indicate that there could be reductions in ICL and ECL when students are presented with DCN. The large fluctuation of this supposed reduction could indicate that the possible benefit of DCN presentation in cognitive load would depend on the learner. ECL in general is high compared to all other surveys, at without visualization and with visualization. If students can adopt new strategies to solve questions using DCN rather than Dirac Notation, a reduction in ICL could be theoretically explainable if these strategies themselves make it easier to solve the task. These results warrant further investigation into the exact conditions under which the presentation of DCN has a positive impact on ICL and ECL.
V.3 Three qubits
In the third survey, participants were asked to find:
the resulting state after application of an X gate, a Z gate, and a Hadamard (H) gate,
the probability of finding a particular result upon measurement of a given state,
the resulting state after a measurement,
the resulting state after application of a CNOT gate,
a fully separable state and a fully entangled state (between partially separable/entangled states), and
a combination of two gates that were applied to transform one given state into another given state (“guess the gates” (Guess-the-Gates) questions).
The average scores for the three-qubit questions are shown in table 8 and the average times for correctly solved questions are shown in table 9. In the latter, the combined score of part A and part B show only cases where participants managed to solve both questions, i.e., two questions of the same topic, correctly. Figure 22 shows the combined average score per question and Figure 23 shows the combined average time taken (for all answers, not just the correct ones). Cases where a participant took more than 10 minutes for one question were excluded.
V.3.1 Three-qubit systems with Circle Notation
In general, the participants scored marginally higher when presented with CN () than when presented the questions without visualization (). The overall score reflects that the questions were suitable for the participants in terms of difficulty. The results of the H gate and Guess-the-Gates questions could indicate that there was a benefit in presenting the visualization to the participants in these contexts, as both groups scored higher here when presented with visualization (0.9 compared to 0.6 for the H gate question and 0.6 compared to 0.1 for the Guess-the-Gates question). In the entanglement question, group math-vis scored higher (0.6 without visualization and 0.4 with visualization) than group vis-math (0.2 with and without visualization) in both parts of the survey. These results of the Entanglement question could indicate differences in context-dependent competence of the two groups, similar to what we observed in the two-qubit survey. In some cases, students seemed to benefit from the visualization, whereas on average there is no difference.
The results warrant further investigation in the regime of the more complex H gate, Guess-the-Gates, Separability, and Entanglement questions to find the reasons for the increases in performance with visualization for some participant groups and large deviations. To benefit from visualization in the context of separability and entanglement, students may need a thorough introduction to the strategy to identify separability and entanglement (see Figure 4). We received feedback from some participants that they did not comprehend the process of identifying entangled states before or during the study. Perhaps, the topic is too advanced as only a subtopic of a study like this and should be introduced more thoroughly in a study with only this context.
Translation competence and cognitive load
In the three-qubit survey, participants presented with CN scored on average on the translation competence questions which is similar to the results of the one- and two-qubit CN survey. Participants who scored perfectly in translation competence also scored comparatively high in the rest of the survey at without visualization and with visualization, compared to and for the other participants. However, we do not see expression of this translation competence in an improvement of score when presented with visualization.
V.3.2 Three-qubit systems with Dimensional Circle Notation
Group math-vis consists of two students and group vis-math of three participants. Hence, the resolution on the average score value per question is rather coarse. Due to the large impact a single person’s performance can have on the results, it is difficult to find any indications of effects. Future studies could focus more on the impact of DCN in the domain of three qubits, in particular. If there is any benefit of visualizing quantum states representing qubits as axes in space, one would expect the benefit to be the largest here because the difference between the dimensional approach and the approach of standard CN is larger the higher the number of qubits in the system. Similarly to the results for the answer correctness, it would be unreasonable to make statements about time taken for correct answers due to the small number of participants.
Translation competence and cognitive load
The impact of the results of one participant seems too large to make any assumptions about the translation competence and cognitive load in the three-qubit DCN context.
V.4 Summary of discussion & Limitations
Considering the suitability of the questions in terms of performance alone and it’s variance between the two groups with or without visualization, we summarize the results of this study in table 1, where we assigned questions a label of easy (average score below 0.51), moderate (between 0.51 and 0.84) or hard (above 0.84), and half the questions with differences between groups and/or whether visualization was presented above the median the high-variance, and below the median the low-variance label. We consider questions of moderate difficulty and questions with higher variance to be of value for further investigation, as it is unclear whether this variance is due to learner characteristics or the presentation of the visualization. Overall, questions of moderate difficulty and high variance in at least two different surveys include the H- and the Z-gate (three-qubit systems), CNOT gate (two-qubit CN and three-qubit DCN), and Entanglement (two qubits). However, in both three-qubit surveys, one participant group answered the Z-gate question perfectly, indicating possible ceiling effects in that question type in particular.
In general, we do not see indications of a large change in performance when participants are presented with a visualization (see table 2), although there are indications of context-dependent performance increases with visualization that require further investigation. This is apparent in the Global Phase Equiv. question of the one-qubit survey, the Z gate, the Measurement Prob., and the separability question of the two-qubit CN survey, the H gate question of the two-qubit DCN survey, and the H gate and Guess-the-Gates questions of the three-qubit CN survey. However, context-dependent learner characteristics seem to have a greater impact on performance than the presence or absence of a visualization. Some groups seem to perform better than other groups specifically in certain types of questions, as we see in the two-qubit CN survey in the X gate, the H gate, separability and entanglement questions, all questions in the two-qubit DCN survey, and in the Z gate, Measurement Prob., entanglement, CNOT gate, and Guess-the-Gates questions in the three-qubit CN survey.
| Question Type | 1-Qubit CN | 2-Qubit CN | 2-Qubit DCN | 3-Qubit CN | 3-Qubit DCN |
| Global Phase Equiv. | m/hv | - | - | - | - |
| X Gate | h | m/hv | e/lv | e/lv | m/lv |
| Z Gate | m | - | - | m/hv | m/hv |
| H Gate | m/lv | h/hv | h/hv | m/hv | m/hv |
| Phase Gate | m/lv | - | - | - | - |
| Measurement Prob. | h/hv | m/lv | h/lv | m/hv | m/lv |
| Measurement State | e/lv | - | - | m/lv | m/hv |
| CNOT Gate | - | m/lv | m/hv | m/lv | m/hv |
| Guess-the-Gates | - | m/lv | h/hv | h/hv | h/lv |
| Separability | - | m/hv | m/lv | h/lv | h/hv |
| Entanglement | - | m/hv | m/hv | h/hv | m/lv |
In order to find indicators of complexity and suitability of questions, we considered the time taken for all answers, not just correct ones (see Figure 16, Figure 21 and Figure 23 for the one-, two- and three-qubit surveys, respectively). During the one-qubit survey, the Global Phase Equiv. and the Hadamard gate question took the participants the longest time. We would omit the Hadamard gate and X gate questions, as they were associated with unnecessary confusion due to incorporation of the global phase equivalency. During the two-qubit survey, the Guess-the-Gates and the Measurement Prob. questions took participants in the CN survey the longest, and the Guess-the-Gates and H gate questions for participants of the DCN survey (see Figure 21). In the three-qubit survey, participants presented with both CN and DCN took the longest in the Guess-the-Gates and H gate questions (see Figure 23), in the former of which some participant groups also performed quite poorly. Due to these results, we would omit the GtG questions in two- and three-qubit systems in future studies. Long times could in some cases be due to greater complexity and in some cases to the distance-difficulty hypothesis [69, 70], which would, in contrast, state that these questions were quite suitable for the participants in terms of level of ability, making the performance in the question an important indicator.
We also measured the cognitive load imposed on the students during the surveys. The generally high ICL and average scores between 0.5 and 0.7 indicate that the difficulty of the problems was adequate overall. We see reductions in cognitive load when students are presented with visualization in both CN and DCN surveys. It could be the case that the amount of space DCN takes up compared to CN leads to increased ECL, due to participants not being able to see all answers at once (which means page scrolling becomes necessary), resulting in a possible split-attention effect [80]. On the other hand, the lower translation competence of participants in the two-qubit DCN survey could also be a reason for the higher ECL. To summarize the results in regard to RQ1, on average, the test items were suitable, even for a heterogeneous target group, with a few exceptions.
Most of the participants were recruited from our quantum computing lecture. While it is an optional Master’s lecture in information science, students from other fields also select it as an elective subject. We expect and have experienced a higher reluctance towards mathematical notation of students from other fields than Physics or Mathematics, especially if not introduced to DN before. When answering the question of whether they have been introduced to DN before, we received the feedback that some students thought that this was before the lecture. We have also received positive feedback regarding the visualizations from the students, while more advanced physics students often do not see a benefit in the visualizations. A possible expertise-reversal effect remains to be investigated. The A-B survey design should cancel out previous knowledge as a relevant factor, if the questions are neither too easy nor too difficult. For future studies with such heterogeneous groups, eye tracking or a think-aloud study design might allow us to gain a better understanding of the underlying thought processes. In studies with more participants, the analysis can be performed by separating the groups according to experience with DN.
In this study, we have seen how performance and cognitive load can be context- and learner-dependent. The results give an overview of the cognitive load with and without visualization, but assumptions cannot be made for specific contexts with these measurements. It can also be taken into account that measurements of cognitive load after multiple questions could be biased toward the most recently answered questions, which would be the measurement questions in the one-qubit survey and the Guess-the-Gates questions in the two- and three-qubit surveys. Performance on these questions is probably very dependent on previous knowledge. This, combined with the high expected heterogeneity, requires separate analyses of the results for different types of questions in future studies.
The test instruments presented in this study are purely designed to measure the benefits of visualization in learning basic quantum operations and entanglement. They cover a broad range of quantum gates in one-, two-, and three-qubit systems, but do not ask of more complicated processes such as whole quantum algorithms, larger qubit-systems and leave out real-world applications like different qubit implementations, discussions about consequences of quantum information theory like the no-cloning theorem, decoherence, and discussion about various applications.
The measures of representational competence presented in this study are limited. The questions measure only a part of representational competence, representational and connectional understanding. Other factors for representational competence would include representational and connectional fluency and meta-representational competencies, considering the symbolic Dirac Notation as a representation and therefore extending the framework in [5]. In the future, additional ideas for questions that provide a more general metric should be generated. For example meta-representational competence of being able to choose between DN, CN, or DCN to most efficiently solve different questions could be investigated using eye tracking. Representational understanding is only measured in the first survey and includes only understanding of the phase in CN. In future studies, this should be extended to finding a measure of representational understanding of also the amplitudes, e.g., by asking students to compare sizes of numbers and sizes of circles. This type of understanding was only included in the translation competence questions.
Due to small sample sizes (especially in the three-qubit DCN survey) and large fluctuations (possibly due to large deviations of learner prerequisites), we can only suspect indications of positive effects of presenting a visualization regarding performance and cognitive load. The circumstances and contexts in which these possible effects occur should be investigated more thoroughly. This study covered a broad range of different contexts, and, therefore, compromises had to be made with regard to the number of questions for each question type. To answer RQ3, reevaluation of the test instruments can be based on fine tuning to make the questions in parts A and B of the surveys more similar. This can be done with respect to questions asking about, e.g., the state after measurement or entanglement, making it so that all participants are presented with amplitude-type or phase-type questions in both parts of the survey. This would either mean an increase in the number of questions or a constriction of question types. Due to the already large number and high complexity of the questions, we recommend the second option.
VI Conclusion & Outlook
In general, the performance and cognitive load results indicate that most of the questions were suitable for the participants, with some notable exceptions of the gate and Measurement State questions possibly being to easy, and the gate, Measurement Probability, Guess-the-Gates, Separability, and Entanglement questions being too hard for some participant groups. Considering the overall time taken, the number of questions is feasible for the participants. The A-B-crossover design allows for within-subject as well between-subject comparisons, while eliminating multiple testing effects due to randomization of the order (A-B or B-A) and we therefore see it as suitable to answer the overarching questions (see Section III).
The results of this study open up a wide range of questions to investigate in future studies. As we see some students benefiting from the visualization (in terms of performance increase and cognitive load reduction) and some not, it will be interesting to investigate further under which circumstances exactly students benefit from the visualization. Especially contexts within the three-qubit DCN setting should be investigated more thoroughly. Here, DCN differs the most from the other two representations, since the number of steps necessary to transform CN to DCN is larger the more qubits the system has. If, in fact, educators wish to teach students about systems of multiple qubits, future studies should also be restricted to systems of multiple qubits. If the goal is teaching only about one-qubit systems, visualizations like the Bloch sphere or vector notation could also be included in the studies.
Due to the observed variances in results, we assume a large learner and context dependency. Therefore, future studies should control for these characteristics in a more detailed way to draw conclusions about the potential benefits of visualizations, as also proposed in [11]. If future studies find significant benefits of visualization, the exact reasons for this could remain unclear, as learners could have used the visualization as support, complementary to the mathematical notation, or have worked exclusively in the visualization. As a solution, eye tracking could be used [81], tracking, for example, the number of transitions between visualization and mathematical symbolism, and the total duration of fixation on visualization and mathematical symbolism. Using eye tracking, it would be possible to see which strategies the participants adopted to solve the questions and which strategies were used by more successful participants. This could result in further design guidelines for educators. It could also be investigated whether there are other benefits of visualization, such as a shorter fixation duration on incorrect answers when presented with visualization [73].
Representational competence as a learner characteristic is theoretically the deciding factor for whether a visualization supports learning [5]. In this study, many students answered these questions perfectly and the number of different representational competence scores was not high. Therefore, it should be considered to measure representational competence more thoroughly in future studies. This could be done, for example, by increasing the number of translation competence questions that measure connectional understanding, and finding other ways to measure representational competence that also includes visual understanding of CN or DCN and quantum operations in these visualizations. Future studies could further investigate the competence of students when presented with only a visualization and how this representational competence impacts the possible benefits of presenting a visualization in addition to DN. Perhaps also, mental rotation ability can help explain possible beneficial effects of DCN [82].
In general, we saw indications of reductions in perceived cognitive load by presenting CN or DCN. Due to the consistency in which this occurs, it is possible that there are beneficial effects regarding a reduction of cognitive load when students are presented with a visualization. A decrease in ECL should, in theory, lead to improved learning outcomes when learning with visualization, if this decrease coincides with an increase in GCL [83, 84]. If we also find a decrease in ICL, it could be that the questions become inherently easier when asked and solved in CN or DCN, which would mean they provide beneficial complementary strategies according to the DeFT framework. Another point to consider is that the cognitive load of learners could be context-dependent, i.e., for some question types, cognitive load difference by presentation of a visualization might differ from other question types. In addition, learning with visualization could also improve the competency of participants when solving questions in uncomplemented DN. Investigating further whether there is a benefit for students presented with visualization, and under which circumstances and in which contexts this occurs, could therefore be of great interest. To control for the high heterogeneity of learners in QIST, this could also be explored using think-aloud interviews [85], or, for studies with higher numbers of participants, separating by experience with DN or quantum physics in general, as we expect higher reluctance towards mathematical notation for those who are not as familiar with it.
Going beyond static multiple external representations, conveying processes in visualizations using video sequences and audio could be worth studying. Furthermore, investigating learning outcomes by including interactivity in the teaching process, for example, by gamification, could be an alternative path forward. There already exists a multitude of quantum games [86] and incorporating CN or DCN into these games is a possibility. Also, the possible beneficial effects of other visualization techniques and their context- and learner-dependency remain to be investigated in multi-qubit systems.
In conclusion, there remains a large space to explore the benefits of visualizations in QIST education. In this work, we presented test instruments to investigate in which circumstances visualizations can support students in the field of QIST and presented first results that narrow down promising contexts for future studies. The measure of representational competence needs rework as there are possible ceiling effects. The other test instruments cover a broad range of contexts and can be adapted and used for further research, also with other visualization methods. However, we recommend more focused studies that use eye tracking and are much more restricted in question types. Restricting the question types will decrease the influence of previous knowledge on the variance of the results. Promising contexts include entanglement, the CNOT gate, and the Hadamard gate in multi-qubit systems, which are also some of the most important concepts of quantum computing. The test items can be designed in ways similar to those described in this study. We expect future studies structured in this way to result in valuable takeaways for educators and see this path forward to be of high interest to the quantum ecosystem.
List of acronyms
- CF
- Competence Framework
- QIST
- Quantum Information Science & Technology
- ECL
- Extraneous Cognitive Load
- GCL
- Germane Cognitive Load
- CN
- Circle Notation
- DN
- Dirac Notation
- DCN
- Dimensional Circle Notation
- ICL
- Intrinsic Cognitive Load
- DeFT
- Design, Functions, Tasks
- MERs
- multiple external representations
- CLT
- Cognitive Load Theory
- STEM
- Science, Technology, Engineering & Mathematics
Appendix A Survey results
The data supporting the findings of this article is openly available [87].
Table 2 summarizes the average translation competence, average scores, and average times for correctly solved questions. In addition, it shows the differences of these performance metrics with visualization minus without visualization. Average ICL, GCL and ECL and the corresponding differences (between with visualization and without visualization) are shown in Table 3. Figure 10 shows the average score per participant, and Figure 11 the average time taken for correctly solved questions. The influence of translation competence on ICL, GCL and ECL as well as the average cognitive load is shown in figures 12, 13 and 14 for tasks with one, two, and three qubits.
| number of qubits | one | two | three | ||
| visualization | CN | CN | DCN | CN | DCN |
| avg. transl. comp. | 0.850.23 | 0.850.14 | 0.710.24 | 0.850.17 | 0.90.12 |
| score w/o vis. | 0.670.24 | 0.620.28 | 0.560.24 | 0.620.26 | 0.690.3 |
| score with vis. | 0.680.24 | 0.630.25 | 0.530.29 | 0.680.25 | 0.620.27 |
| score difference | 0.00.29 | 0.010.22 | -0.020.27 | 0.060.1 | -0.070.15 |
| time w/o vis. | 10474 | 8247 | 12158 | 10532 | 148136 |
| time with vis. | 8856 | 11376 | 13999 | 11976 | 11661 |
| time difference | -1665 | 3251 | 1767 | 1457 | -3287 |
| number of qubits | one | two | three | ||
| visualization | CN | CN | DCN | CN | DCN |
| ICL w/o vis. | 3.750.97 | 4.20.6 | 4.71.49 | 4.620.99 | 4.41.36 |
| ICL with vis. | 3.641.04 | 3.890.99 | 4.01.54 | 4.41.11 | 4.01.0 |
| ICL difference | -0.081.07 | -0.220.79 | -0.51.28 | -0.251.2 | 0.01.22 |
| GCL w/o vis. | 4.121.41 | 4.670.82 | 4.911.0 | 4.891.37 | 4.61.5 |
| GCL with vis. | 3.941.26 | 4.110.74 | 4.731.05 | 4.41.36 | 4.01.41 |
| GCL difference | -0.121.27 | -0.620.99 | -0.180.83 | -0.441.26 | -0.61.62 |
| ECL w/o vis. | 2.881.67 | 2.670.75 | 3.951.42 | 3.051.17 | 3.31.6 |
| ECL with vis. | 2.761.54 | 2.720.82 | 3.641.62 | 2.751.33 | 2.51.1 |
| ECL difference | -0.122.21 | 0.061.34 | -0.321.5 | -0.31.33 | -0.82.11 |
In the following, we show detailed results for single-qubit (Section A.1), two-qubit (Section A.2) and three-qubit studies (Section A.3). We describe the average correctness per question with and without visualization and survey parts A and B and the combined score of part A and B. Then, we do the same for average time taken for correct answers, considering for the combination of parts A and B only cases where both questions of each topic, with and without visualization, were answered correctly.
A.1 One-qubit survey
The average scores for single-qubit questions are shown in table 4 and average times for correctly solved questions in table 5. In the latter, the combined score of part A and part B show only cases where participants managed to solve both questions, i.e. both questions on the same concept with and without visualization, correctly. Figure 15 shows the combined average score per question and Figure 16 shows the combined average time taken (for all answers, not just the correct ones). Cases in which a participant took more than 20 minutes for one question were excluded.
During part A, group math-vis solved questions without visualization with an average score of and an average time of minutes for correctly solved questions, while group vis-math answered the same questions with visualization with an average score of and an average time of minutes. During part B, the average scores were 0.570.17 (group vis-math) and (group math-vis), respectively, at average times for correctly solved questions of and minutes. When combining parts A and B of the survey, the average scores were without visualization and with visualization, at and minutes.
In survey part A, group math-vis scored higher than group vis-math on the global phase, X gate, phase gate, H gate, Measurement Prob., and Measurement State questions with an average score of to on these questions and did not score lower on any questions. Group math-vis took longer on correct answers to the X gate, phase gate, H gate, and Measurement State questions with an average of to seconds while taking less time on the Global Phase Equiv. and Measurement Prob. questions with to seconds.
In part B of the survey, group math-vis scored higher than group vis-math on the Global Phase Equiv., Z gate, H gate, Measurement Prob., and Measurement State questions, with an average of to . Group vis-math scored higher on the phase gate question, with 0.88 to 0.78. Group math-vis took more time on the global phase, Z gate, and Measurement State questions with to seconds and group vis-math took longer on phase gate, H gate and Measurement Prob. questions with to seconds.
When presented with visualization, group math-vis scored higher on the global phase, H gate, Measurement Prob., and Measurement State questions with average scores of to . When presented questions with visualization, group math-vis took longer on the Global Phase Equiv. question with 177 to 77 seconds and took a shorter amount of time on the phase gate, H gate, Measurement Prob., and Measurement State questions with to seconds than when the group was not presented with a visualization. Group vis-math scored higher when presented without visualization on the global phase, phase gate, H gate, and Measurement Prob. question with scores of to . Group vis-math took longer on the Measurement Prob. and Measurement State questions when presented with visualization with to seconds and took less time on the global phase, phase gate, and H gate questions with to seconds.
| survey part | A | B | combined | |||
| question | w/o vis.111N = 9 | with vis.222N = 8 | w/o vis.222N = 8 | with vis.111N = 9 | w/o vis.333N = 17 | with vis.333N = 17 |
| Global Phase Equiv. | 0.56 | 0.5 | 0.62 | 1.0 | 0.59 | 0.76 |
| X gate | 0.56 | 0.25 | - | - | - | - |
| Z gate | - | - | 0.62 | 0.78 | - | - |
| phase gate | 0.78 | 0.75 | 0.88 | 0.78 | 0.82 | 0.76 |
| H gate | 0.67 | 0.62 | 0.75 | 0.78 | 0.71 | 0.71 |
| Measurement Prob. | 0.56 | 0.12 | 0.38 | 0.67 | 0.47 | 0.41 |
| Measurement State | 0.89 | 0.75 | 0.75 | 1.0 | 0.82 | 0.88 |
| Mean | 0.670.13 | 0.50.24 | 0.670.16 | 0.830.12 | 0.680.15 | 0.710.18 |
| survey part | A | B | combined | |||
| question | w/o vis.111N = 9 | with vis.222N = 8 | w/o vis.222N = 8 | with vis.111N = 9 | w/o vis.333N = 17 | with vis.333N = 17 |
| Global Phase Equiv. | 1.280.63(6) | 1.481.57(5) | 2.942.99(6) | 2.952.2(10) | 1.631.08(10) | 2.262.27(10) |
| X gate | 4.312.61(5) | 1.520.92(3) | - | - | - | - |
| Z gate | - | - | 3.695.56(5) | 5.434.11(7) | - | - |
| phase gate | 1.652.49(8) | 1.010.44(7) | 1.191.05(9) | 0.870.42(7) | 1.471.99(11) | 0.930.46(11) |
| H gate | 2.932.9(7) | 0.680.41(6) | 1.782.74(6) | 1.50.83(7) | 2.132.59(10) | 0.840.4(10) |
| Measurement Prob. | 1.231.0(6) | 1.33(1) | 0.860.59(4) | 0.810.58(6) | 1.130.96(6) | 0.910.61(6) |
| Measurement State | 1.141.39(9) | 0.550.24(7) | 0.380.17(6) | 0.540.68(9) | 0.831.14(14) | 0.550.55(14) |
| Mean | 2.091.27 | 1.090.41 | 1.811.28 | 2.011.88 | 1.440.49 | 1.10.67 |
A.1.1 Representational competence
Participants were asked to calculate Euler formula identities and similar questions to translate between CN to a complex number. In addition, they were asked to multiply numbers written in DN and CN as shown in Figure 7. Then they were asked to rate their confidence in the question from 0 (random guess), 1 (very unsure) to 4 (very sure). The correct answers were weighted (multiplied by , where is the confidence). The weighted (unweighted) average mathematics representational understanding was (). The weighted (unweighted) average CN representational understanding was (). The average weighted (unweighted) difference of CN representational understanding minus mathematics representational understanding was ().
Figure 17 shows, for each participant in the first survey, the difference in scores (with visualization minus without visualization) and Figure 18 the difference in time, both plotted against the difference in representational understanding between visualization and math. Figure 19 shows the dependency of the difference in intrinsic, germane, and ECL on the representational understanding difference.
A.1.2 Translation competence and cognitive load
On average translation competence, 12 students achieved perfect scores and 5 students achieved a score of 0.5. The students who did not score perfectly also scored lower in the rest of the survey ( to without visualization and to with visualization) and took a smaller amount of time on the correctly solved questions ( to seconds without visualization and to seconds with visualization) than the students who solved the translation questions perfectly.
A.2 Two-qubit survey
Table 6 lists the average scores and table 7 lists the average time taken for correctly solved one-qubit questions, with and without visualization, for surveys that were taken supported by CN and DCN. Figure 20 shows the combined average score per question and Figure 21 shows the combined average time taken (for all answers, not just correct ones). Cases where a participant took more than 20 minutes for one question were excluded.
| Circle Notation | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.111N = 10 | with vis.111N = 10 |
| X gate | 1.0 | 0.4 | 0.6 | 0.8 | 0.8 | 0.6 |
| H gate | 0.8 | 0.2 | 0.4 | 0.6 | 0.6 | 0.4 |
| Measurement Prob. | 0.6 | 0.6 | 0.4 | 0.6 | 0.5 | 0.6 |
| Separability | 0.6 | 1.0 | 0.8 | 0.6 | 0.7 | 0.8 |
| Entanglement | 0.8 | 1.0 | 0.6 | 0.4 | 0.7 | 0.7 |
| CNOT gate | 0.8 | 0.8 | 0.8 | 0.6 | 0.8 | 0.7 |
| Guess-the-Gates | 0.6 | 0.6 | 0.4 | 0.6 | 0.5 | 0.6 |
| Mean | 0.740.14 | 0.660.28 | 0.570.17 | 0.60.11 | 0.660.12 | 0.630.12 |
| Dimensional Circle Notation | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.333N = 6 | with vis.222N = 5 | w/o vis.222N = 5 | with vis.333N = 6 | w/o vis.444Standard deviation is used for confidence intervals.Confidence intervals are the average of the standard deviations. | with vis.444N = 3 |
| X gate | 1.0 | 0.8 | 0.8 | 0.83 | 0.91 | 0.82 |
| H gate | 0.17 | 0.6 | 0.4 | 0.17 | 0.27 | 0.36 |
| Measurement Prob. | 0.5 | 0.6 | 0.6 | 0.33 | 0.55 | 0.45 |
| Separability | 0.5 | 0.4 | 0.6 | 0.83 | 0.55 | 0.64 |
| Entanglement | 0.5 | 0.8 | 0.6 | 0.33 | 0.55 | 0.55 |
| CNOT gate | 0.67 | 1.0 | 0.8 | 0.5 | 0.73 | 0.73 |
| Guess-the-Gates | 0.67 | 0.2 | 0.0 | 0.17 | 0.36 | 0.18 |
| Mean | 0.570.23 | 0.630.25 | 0.540.26 | 0.450.26 | 0.560.2 | 0.530.2 |
| Circle Notation111N = 11 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.222N = 6 | with vis.333N = 11 | w/o vis.333N = 5 | with vis.222N = 6 | w/o vis.111N = 11 | with vis.111N = 11 |
| X gate | 0.710.16(5) | 1.340.06(2) | 2.051.06(3) | 0.80.35(5) | 0.950.39(6) | 1.080.23(6) |
| H gate | 0.930.96(5) | 2.18(1) | 2.280.83(2) | 2.20.64(3) | 1.150.96(3) | 2.20.64(3) |
| Measurement Prob. | 2.380.33(3) | 4.582.83(3) | 1.380.34(2) | 4.124.42(3) | 2.00.67(4) | 4.393.42(4) |
| Separability | 1.781.34(3) | 1.870.61(5) | 0.950.94(4) | 0.840.25(3) | 1.31.2(7) | 1.380.7(7) |
| Entanglement | 1.470.39(4) | 1.731.21(5) | 1.090.33(3) | 1.240.12(2) | 1.360.43(5) | 1.861.13(5) |
| CNOT gate | 1.591.77(5) | 2.141.31(4) | 0.920.57(4) | 2.080.77(3) | 1.651.41(7) | 2.111.11(7) |
| Guess-the-Gates | 2.342.01(3) | 2.080.61(3) | 2.342.29(2) | 1.171.04(3) | 2.412.37(4) | 1.480.91(4) |
| Mean | 1.60.64 | 2.281.06 | 1.570.63 | 1.781.17 | 1.550.51 | 2.071.1 |
| Dimensional Circle Notation111N = 11 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.222N = 6 | with vis.333N = 5 | w/o vis.333N = 5 | with vis.222N = 6 | w/o vis.111N = 11 | with vis.111N = 11 |
| X gate | 1.130.38(6) | 1.770.67(4) | 1.350.65(4) | 1.70.66(5) | 1.290.49(9) | 1.730.67(9) |
| H gate | 2.73(1) | 6.127.98(3) | 6.171.65(2) | 5.88(1) | 7.82(1) | 17.4(1) |
| Measurement Prob. | 3.221.7(3) | 2.911.62(3) | 1.710.26(3) | 5.133.27(2) | 1.860.34(4) | 2.651.48(4) |
| Separability | 2.431.85(3) | 3.081.99(2) | 2.482.22(3) | 1.680.9(5) | 2.832.03(5) | 2.141.51(5) |
| Entanglement | 2.270.49(3) | 1.391.4(4) | 1.410.95(3) | 2.772.48(2) | 1.810.95(5) | 2.092.04(5) |
| CNOT gate | 2.451.94(4) | 1.571.0(5) | 1.560.87(4) | 1.770.75(3) | 2.11.65(7) | 1.830.83(7) |
| Guess-the-Gates | 2.530.68(4) | 1.9(1) | (0) | 7.07(1) | (0) | (0) |
| Mean | 2.390.64 | 2.681.65 | 2.441.87 | 3.712.27 | 2.952.44 | 4.646.26 |
A.2.1 Circle Notation
During part A, participants assigned to the CN group solved the questions without visualization with an average score of with an average time of minutes for correctly solved questions, while the other group answered the same questions with visualization with an average score of with an average time of minutes. During part B, the mean scores with and without visualization were and , respectively, at average times for correctly solved questions of and minutes. When combining parts A and B of the survey, the average scores were without CN and with CN, at and minutes.
While group math-vis, that was assigned questions without CN first, scored higher than group vis-math on the X gate and H gate questions in survey part A with, on average, to , the inverse is apparent for the Separability and Entanglement questions, where group vis-math scored higher than group math-vis 1.0 to . Group vis-math, with CN, took longer to correctly solve the X gate, H gate, Measurement Prob., Separability, Entanglement, and CNOT gate questions with to seconds and took less time on the Guess-the-Gates questions with 140 to 125 seconds.
In survey part B, group math-vis (that was now given questions with CN) scored higher than group vis-math on the X gate, H gate, Measurement Prob., and Guess-the-Gates questions with respective average scores to and group vis-math scored higher than group math-vis on the separability, entanglement, and CNOT gate questions with to . In this survey part, group math-vis, with CN, took longer on the Measurement Prob., entanglement, and CNOT gate questions with, on average, to seconds while taking less time on the X gate, H gate, separability, and Guess-the-Gates questions with to seconds on these questions on average.
Comparing within each group, group math-vis scored worse with CN than without CN on the X gate, H gate, entanglement, and CNOT gate questions with to . When solving questions correctly, group math-vis took longer on the X gate, H gate, Measurement Prob., and CNOT gate questions when presented with CN, with to seconds, while taking a shorter amount of time on the separability, entanglement, and Guess-the-Gates questions with to seconds.
When presented with CN, group vis-math scored higher than when not presented with a visualization on the Measurement Prob., separability, entanglement, and Guess-the-Gates questions with, on average, with to . The group scored higher when solving questions without CN on the X gate and H gate questions with to . They took longer with CN when solving the Measurement Prob., separability, entanglement, and CNOT gate questions correctly, with an average of to seconds, and took less time on the X gate, H gate, and Guess-the-Gates questions when presented with CN, with to seconds.
A.2.2 Dimensional Circle Notation
During part A, participants assigned to the DCN group solved questions without visualization with an average score of with an average time of minutes for correctly solved questions, while the same questions with visualization were answered with an average score of with an average time of minutes. During part B, the average scores were and , respectively, at average times for correctly solved questions of and minutes. When combining the survey parts A and B, the average scores were without visualization and with visualization, at and minutes.
In both survey parts, on the X gate, Separability, and Guess-the-Gates questions, group math-vis scored higher than group vis-math. In part A, this was to and in part B, with to . On the other hand, group vis-math scored higher than group math-vis on the H gate, Measurement Prob., Entanglement, and CNOT gate questions to in part A and to in part B.
In survey part A, group math-vis took longer on the H gate, Measurement Prob., Entanglement, CNOT gate, and Guess-the-Gates questions with to seconds and less time on the X gate and Separability questions with to seconds. In survey part B, group math-vis took longer on the X gate, Measurement Prob., Entanglement, and CNOT gate questions with to seconds and group vis-math took longer on the H gate and Separability questions with to seconds.
Group math-vis had a higher score on the separability questions with DCN with to while performing worse with DCN on the X gate, Measurement Prob., Entanglement, CNOT, and Guess-the-Gates questions with to . The group took longer with DCN when correctly solving the X gate, H gate, Measurement Prob., Entanglement, and Guess-the-Gates questions with to seconds and took less time on the Separability and CNOT questions with to seconds.
Supported by DCN, group vis-math scored higher on the H gate, Entanglement, CNOT, and Guess-the-Gates questions with to while scoring worse on the Separability question with to . On the X gate, Measurement Prob., Separability, and CNOT questions, group vis-math took longer with DCN, with to seconds, while taking longer without DCN on the H gate and Entanglement questions with to seconds.
A.3 Three-qubit survey
Table 8 lists the average scores and table 9 lists the average time taken for correctly solved one-qubit questions, with and without visualization, for surveys that were taken supported by CN and DCN. Figure 22 shows the combined average score per question and Figure 23 shows the combined average time taken (for all answers, not just the correct ones). Cases in which a participant took more than 20 minutes for one question were excluded.
| Circle Notation111N = 10 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.111N = 10 | with vis.111N = 10 |
| X gate | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
| Z gate | 0.6 | 1.0 | 1.0 | 0.6 | 0.8 | 0.8 |
| H gate | 0.6 | 0.8 | 0.6 | 1.0 | 0.6 | 0.9 |
| Measurement Prob. | 0.6 | 1.0 | 1.0 | 0.6 | 0.8 | 0.8 |
| Measurement State | 0.6 | 0.6 | 0.8 | 0.6 | 0.7 | 0.6 |
| Separability | 0.4 | 0.4 | 0.4 | 0.2 | 0.4 | 0.3 |
| Entanglement | 0.6 | 0.2 | 0.2 | 0.4 | 0.4 | 0.3 |
| CNOT gate | 0.8 | 1.0 | 0.8 | 0.6 | 0.8 | 0.8 |
| Guess-the-Gates | 0.0 | 0.8 | 0.2 | 0.4 | 0.1 | 0.6 |
| Mean | 0.580.26 | 0.760.28 | 0.670.31 | 0.60.25 | 0.620.28 | 0.680.25 |
| Dimensional Circle Notation222N = 5 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.333N = 2 | with vis.444N = 3 | w/o vis.444N = 3 | with vis.333N = 2 | w/o vis.222N = 5 | with vis.222N = 5 |
| X gate | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| Z gate | 1.0 | 0.67 | 0.33 | 1.0 | 0.6 | 0.8 |
| H gate | 1.0 | 0.0 | 0.67 | 0.5 | 0.8 | 0.2 |
| Measurement Prob. | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| Measurement State | 0.5 | 1.0 | 0.33 | 1.0 | 0.4 | 1.0 |
| Separability | 1.0 | 0.33 | 0.67 | 0.0 | 0.8 | 0.2 |
| Entanglement | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| CNOT gate | 1.0 | 0.67 | 0.33 | 0.5 | 0.6 | 0.6 |
| Guess-the-Gates | 0.5 | 0.33 | 0.67 | 0.5 | 0.6 | 0.4 |
| Mean | 0.890.21 | 0.560.27 | 0.560.16 | 0.720.34 | 0.690.15 | 0.620.29 |
| Circle Notation111N = 10 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.222N = 5 | with vis.222N = 5 | w/o vis.111N = 10 | with vis.111N = 10 |
| X gate | 0.940.36(5) | 1.461.16(5) | 1.520.82(5) | 1.050.34(5) | 1.230.7(10) | 1.260.88(10) |
| Z gate | 1.030.04(3) | 2.021.18(5) | 1.530.87(5) | 1.170.29(3) | 1.350.73(8) | 1.71.04(8) |
| H gate | 3.721.74(3) | 6.825.99(4) | 3.641.59(3) | 1.850.53(5) | 3.681.67(6) | 4.795.26(6) |
| Measurement Prob. | 1.420.66(3) | 2.71.17(5) | 1.270.87(5) | 0.440.2(3) | 1.320.8(8) | 1.851.43(8) |
| Measurement State | 1.920.1(3) | 3.531.2(3) | 2.751.71(4) | 1.280.44(3) | 2.141.31(6) | 2.411.44(6) |
| Separability | 1.190.36(2) | 3.11.92(2) | 2.180.28(2) | 1.27(1) | 2.010.46(2) | 1.230.04(2) |
| Entanglement | 2.532.24(3) | 2.1(1) | 2.27(1) | 1.160.14(2) | 3.071.89(3) | 1.470.46(3) |
| CNOT gate | 1.410.46(4) | 1.681.72(5) | 1.60.94(4) | 1.060.04(3) | 1.530.8(7) | 1.481.49(7) |
| Guess-the-Gates | (0) | 3.591.35(4) | 5.13(1) | 1.980.42(2) | 5.13(1) | 5.13(1) |
| Mean | 1.770.94 | 3.01.63 | 2.431.26 | 1.250.45 | 2.381.33 | 2.371.51 |
| Dimensional Circle Notation222N = 5 | ||||||
| survey part | A | B | combined | |||
| question | w/o vis.333N = 2 | with vis.444N = 3 | w/o vis.444N = 3 | with vis.333N = 2 | w/o vis.222N = 5 | with vis.222N = 5 |
| X gate | 1.830.42(2) | 3.112.38(2) | 1.490.43(2) | 1.940.69(2) | 1.660.45(4) | 2.521.84(4) |
| Z gate | 1.230.55(2) | 0.970.1(2) | 1.95(1) | 1.881.25(2) | 1.470.56(3) | 1.611.09(3) |
| H gate | 8.023.47(2) | (0) | 1.790.51(2) | 1.32(1) | 4.55(1) | 1.32(1) |
| Measurement Prob. | 2.431.0(2) | 1.120.04(2) | 0.520.09(2) | 2.351.9(2) | 1.481.19(4) | 1.741.48(4) |
| Measurement State | 1.05(1) | 1.080.2(3) | 1.77(1) | 2.01.4(2) | 1.410.36(2) | 0.830.23(2) |
| Separability | 3.662.66(2) | 1.57(1) | 3.513.28(2) | (0) | 6.78(1) | 1.57(1) |
| Entanglement | 6.124.09(2) | 1.880.52(2) | 2.11.93(2) | 2.761.43(2) | 4.113.78(4) | 2.321.16(4) |
| CNOT gate | 0.890.18(2) | 1.550.62(2) | 2.68(1) | 4.52(1) | 1.880.81(2) | 2.721.79(2) |
| Guess-the-Gates | 17.48(1) | 7.28(1) | 2.181.78(2) | 1.52(1) | 3.97(1) | 7.28(1) |
| Mean | 4.755.37 | 2.322.12 | 2.00.82 | 2.291.01 | 3.031.91 | 2.441.92 |
A.3.1 Circle Notation
During part A, participants assigned to the CN group solved questions without visualization with an average score of and an average time of minutes for correctly solved questions, while the same questions with visualization were answered with an average score of and an average time of minutes. During part B, the average scores were and , respectively, at average times for correctly solved questions of and minutes. When combining parts A and B of the survey, the average scores were without visualization and with visualization, at and minutes.
In survey part A, group math-vis, that was presented questions without CN first, scored higher than group vis-math on the Entanglement question with 0.6 to 0.2. Group vis-math scored higher on the Z gate, the H gate, the Measurement Prob., the CNOT, and the Guess-the-Gates questions with to . In this survey part, group math-vis took longer when solving the Entanglement question correctly, with 152 to 126 seconds, while group vis-math took longer for correct solutions to the X gate, Z gate, H gate, Measurement Prob., Measurement State, Separability, and CNOT gate questions, with to seconds.
In survey part B, group math-vis scored higher on the H gate, entanglement, and Guess-the-Gates questions with to , while group vis-math scored higher on the Z gate, Measurement Prob., Measurement State, Separability, and CNOT gate questions with to . Group vis-math took longer on all questions in survey part B with to seconds.
Group math-vis, that was presented questions without CN first, scored higher on the H gate and Guess-the-Gates questions with to with CN. The group scored worse with CN on the separability, entanglement, and CNOT gate questions by a score difference of to. With CN. Participants in the group took longer when solving the X gate, Z gate and Separability questions correctly, to seconds, while taking less time on the H gate, Measurement Prob., Measurement State, Entanglement, CNOT gate, and Guess-the-Gates questions with to seconds.
Group vis-math, that was presented questions with CN first, scored higher on the H gate, CNOT gate, and Guess-the-Gates questions with CN with to . The group scored lower with CN on the Measurement State question with 0.6 as compared to 0.8. When presented with CN, participants in the group took longer on the Z gate, H gate, Measurement Prob., Measurement State, Separability, and CNOT gate questions, with to . They took less time on the X gate, entanglement, and Guess-the-Gates questions, with to seconds.
A.3.2 Dimensional Circle Notation
During part A, participants assigned to the DCN group solved questions without visualization with an average score of with an average time of minutes for correctly solved questions, while the same questions with visualization were answered with an average score of 0.60.24 with an average time of minutes. During part B, the average scores were and , respectively, at average times for correctly solved questions of and minutes. When combining parts A and B of the survey, the average scores were without visualization and with visualization, at and minutes.
Group math-vis, that was assigned questions without DCN first, scored higher than group vis-math on the X gate, Z gate, H gate, Measurement Prob., separability, entanglement, CNOT, and Guess-the-Gates questions in survey part A, with to . The inverse is apparent for the Measurement State question, where group vis-math scored higher than group math-vis with 1 to 0.5. Group vis-math, with DCN, took longer to correctly solve the Z gate, Measurement Prob., Separability, Entanglement, and Guess-the-Gates questions, with to seconds, whereas group math-vis took longer for correctly solving the X gate, Measurement State, and CNOT gate questions, to seconds.
In survey part B, group math-vis (that now was presented questions with DCN) scored higher than group vis-math on the X gate, Z gate, Measurement Prob., Measurement State, Entanglement, and CNOT gate questions with to and group vis-math scored higher than group math-vis on the H gate, separability, and Guess-the-Gates questions with to . In this survey part, group math-vis, with DCN, took longer on the X gate, Measurement Prob., Measurement State, Entanglement, and CNOT gate questions, with to seconds, while taking less time on the Z gate, H gate, and Guess-the-Gates questions, with to seconds.
Comparing within each group, group math-vis scored higher with DCN on the Measurement State question with 1 to 0.5, and scored lower with DCN than without DCN on the H gate, Separability, and CNOT gate questions by averages of to . When solving questions correctly, group math-vis took longer on the X gate, Z gate, Measurement State, and CNOT gate questions when presented with DCN, with to seconds, while taking less time on the H gate, Measurement Prob., Entanglement, and Guess-the-Gates questions, to seconds.
When presented with DCN, group vis-math scored higher than itself on the Z gate, Measurement State, and CNOT gate questions by, on average, to . The group scored higher when solving questions without DCN on the H gate, Separability, and Guess-the-Gates questions with to . They took longer with DCN when solving the X gate, Measurement Prob., and Guess-the-Gates questions correctly, to seconds, and took less time on the Z gate, Measurement State, separability, Entanglement, and CNOT gate questions when presented with DCN, on average to seconds.